A solution of carbon tetrachloride and acetic acid CH3COOH is boiling at 107.6°C. A sample of the vapor above the solution is cooled until it condenses. This condensed sample is analyzed, and turns out to be 50.% carbon tetrachloride by mass.Calculate the percent by mass of carbon tetrachloride in the boiling solution. Here's some data you may need:

normal boiling point density vapor pressure at

107.6°C

carbon tetrachloride

77.°C

1.6gmL

1733.torr

acetic acid

118.°C

1.1gmL

623.torr

Be sure your answer has the correct number of significant digits.

Note for advanced students: you may assume the solution and vapor above it are ideal.
Answer #1

Suppose Vabours maiss - PISS = 100 g If caly then CH₃COOH ll ll = 509 ll Nechy a 5/154 0-324 Notal o Xcely 0•324 1 Total 50/1and X8 = 1-00122=C0-878) Noce we have male fraction one 1 1 1 1 1 From this we can calculete mass rose calculation > of cely